OpenStudy (anonymous):

please help me with Algebra 1

OpenStudy (anonymous):

1. What is the simplified form of the picture below? a. x3 – 3x2 + 9x – 27 b. (x2 + 9)(x2 – 9) c. x3 + 3x2 + 9x + 27 d. x3 + 3x2 – 9x – 27

OpenStudy (anonymous):

a^2-b^2= (a+b)(a-b)

OpenStudy (anonymous):

factorise x^4-81 using the above identity

OpenStudy (anonymous):

I still do not understand. I suck at algebra lol

OpenStudy (anonymous):

What is \[(x^2)^2\] ?

OpenStudy (anonymous):

umm x^4 ?

OpenStudy (anonymous):

What is square root of 81?

OpenStudy (anonymous):

9?

OpenStudy (anonymous):

So x^4-81=(x^2)^2-9^2 Then factorise this

OpenStudy (anonymous):

is it c?